Inequalities ; If xy=25 find the least possible value for x+y, if both x and y are positive?
Answer by stanbon(75887)

If xy=25 find the least possible value for x+y, if both x and y are positive?
---
x = 25/y
---
Substitute to get:
Sum = (25/y)+y
---
Sum = (25+y^2)/y
----
Find the derivative:
S' = [y(2y)-(25+y^2)]/y^2
--
S' =[y^2-25]/y^2
---
S' = 1-[(25)/y^2]
----
To find min/max solve S' = 0
1-[(25/y^2)] = 0
(25/y^2) = 1
y^2 = 25
y = 5
---
Substitute xy = 25
If y = 5, then x = 5
---
So, the minimum sum is 5+5 = 10
==================================
Cheers,
Stan H.